OpenStudy (anonymous):

Given the functions f(x) = x(x2 - 3) and g(x) = 3 + 2x, complete the table for each function.

OpenStudy (lynfran):

let x= various values example x=0,1,2,3,5,8,etc to find the function's corresponding y values example when x=0 f(0)=0(0^2-3)..=0 g(0)=3+2(0)..3 so when x=0 , f(0)=0 and g(0)=3 ...follow same format

OpenStudy (lynfran):

or did they give a specific table to be completed?